Circular No. 7465

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 O1 (KOEHN) MPEC 2000-O22 shows the identification, by B. G. Marsden, of this

comet with the LINEAR “asteroidal” objects 1998 XA_{70} and 1999 UJ_{10}, together with improved orbital elements: T = 2000 Jan. 27.35 TT, ω = 55^{o}.11, Ω = 88^{o}.86, i = 148^{o}.10 (equinox 2000.0), q = 5.9219 AU, e = 1.0009.

HD 155826 C. Lisse and A. Schultz, Space Telescope Science Institute; Y. Fer-

nandez, University of Hawaii; S. Peschke, European Space Agency; M. Ressler, Jet Propulsion Laboratory, NASA; D. Backman, Ames Research Center, NASA; and C. Kaminski and B. Golisch, Infrared Telescope Facility (IRTF), NASA, report the discovery of an extremely red object located 7′′ northwest (p.a. 323^{o}) of the bright binary system HD 155826: “The infrared cameras NSFCAM and MIRLIN were used at the IRTF on the nights of June 18 and 19 to image the area around HD 155826. Originally reported by IRAS as one source detected at 12–100 μm, within the error ellipse of HD 155826, two bright point sources were found in MIRLIN’s broadband N field at 0.63 and 1.3 Jy. HD 155826 is easily detected in our NSFCAM z, J, H, and K′ images, but there is no apparent counterpart for the new source down to mag 14 in Gunn z, J, or the Aladin sky catalogue. The new source may have been detected, however, at low S/N by NSFCAM in H and K′ at mag 14 and 13, respectively. The detection of the second, non-optical, source explains the confusing K–(IRAS 12μm) colors originally assigned to HD 155826 as a possible Vega-like system (Aumann 1985, PASP 97, 885; Walker and Wolstencroft 1988, PASP 100, 1509; Sylvester et al. 1996, MNRAS 279, 915). Assuming that the new object and HD 155826 are associated, the 1′′.3 FWHM point-spread-function diameter at 10 μm implies that any unresolved source at the distance of HD 155826, 31 pc, should be ≤ 33 AU in size. A fit to the IRAS long-wavelength fluxes yields a color temperature of ∼ 130 K and a lower size limit for an optically thick source with 1.3 Jy at 10 μm of ∼ 2 AU. The field is within 12^{o} of the Galactic center and < 7′ from the Galactic plane, so the possibility of a chance alignment of two unconnected systems cannot yet be ruled out. Barring that, we conclude that the most likely identity of the ‘companion’ is either a protostar or carbon star with a thick shell (it is too small and hot to be cirrus, too big to be a planet, and too faint at J and K to be a brown dwarf).”

2000 July 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green